OpenStudy (anonymous):

Given that f(x) = x2 − 3x + 3 and g(x) = the quantity of x minus one, over four, solve for f(g(x)) when x = 5. −3 1 7 13

jimthompson5910 (jim_thompson5910):

g(5) = ???

OpenStudy (anonymous):

idk what g =

OpenStudy (anonymous):

idk what g equels

jimthompson5910 (jim_thompson5910):

the function g(x) is given to you

jimthompson5910 (jim_thompson5910):

you just replace every x in g(x) with 5 and evaluate

OpenStudy (anonymous):

ok i did that and i got -8

jimthompson5910 (jim_thompson5910):

you got g(5) = -8 ?

OpenStudy (anonymous):

oh nevermind its 1

OpenStudy (anonymous):

because 4/4 = 1 correct?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

then you plug that into f(x)

OpenStudy (anonymous):

ok i got 4

jimthompson5910 (jim_thompson5910):

incorrect

OpenStudy (anonymous):

i thought so because that wasnt one of the answer choices

OpenStudy (anonymous):

come on jim dont igve up on me

jimthompson5910 (jim_thompson5910):

f(x) = x^2 − 3x + 3 f(1) = (1)^2 − 3(1) + 3 f(1) = ???

OpenStudy (anonymous):

what does that symbole mean in between the x and 2?

OpenStudy (anonymous):

i get -4

jimthompson5910 (jim_thompson5910):

^ means exponent

jimthompson5910 (jim_thompson5910):

x^2 means "x squared"

OpenStudy (anonymous):

first solve for g(5) and then whatever you get as your answer for that use as f(?)

OpenStudy (anonymous):

ohhh ok yeah i got it now the answer is -3

jimthompson5910 (jim_thompson5910):

(1)^2 = 1*1 = 1 other examples 3^2 = 3*3 = 9 5^2 = 5*5 = 25 etc etc

OpenStudy (anonymous):

thanks alot man

jimthompson5910 (jim_thompson5910):

the final answer isn't -3

OpenStudy (anonymous):

you said you got 1 as your answer for g(5). now solve for f(1). to do that take the function you were given for f and substitute 1 in for every x you see and solve

OpenStudy (anonymous):

its like @jim_thompson5910 said earlier. when they wrote "f(x) = x^2 − 3x + 3 f(1) = (1)^2 − 3(1) + 3 f(1) = ???"

OpenStudy (anonymous):

try just solving the second line they wrote up there

OpenStudy (anonymous):

thats what i did and i get -3

OpenStudy (anonymous):

1^2 = ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1 - 3

OpenStudy (anonymous):

then +3

OpenStudy (anonymous):

or just 1 - 0 XD

OpenStudy (anonymous):

oh ok so the answer is 1

OpenStudy (anonymous):

yep awesome

OpenStudy (anonymous):

thanks for your time !

OpenStudy (anonymous):

haha no problem. happy to help